Marshall Inc. recently hired your consulting firm to improve the company's performance. It has been highly profitable but has been experiencing cash shortages due to its high growth rate. As one part of your analysis, you want to determine the firm's cash conversion cycle. Using the following information and a 365-day year, what is the firm's present cash conversion cycle?

Average inventory = $75,000
Annual sales = $600,000
Annual cost of goods sold = $360,000
Average accounts receivable = $180,000
Average accounts payable = $37,000
Answer #1
Cash conversion cycle = DIO+DSO-DPO
DIO = Days inventory outstanding = Average inventory*365/COGS = 75000*365/360000 = 76.04 Days
DSO = Days sales outstanding = Average accounts receivable*365/Total credit sales = 180000*365/600000 = 109.50 Days
DPO = Days payable outstanding = Average accounts payable*365/COGS = 37000*365/360000 = 37.51 Days
Cash conversion cycle = DIO+DSO-DPO = 76.04+109.50-37.51 = 148.03 Days